OpenStudy (anonymous):

The average of two numbers is equal to twice the positive difference between the two numbers. If the larger number is 35, what is the smallest number?

Parth (parthkohli):

Can you introduce suitable notation here?

Parth (parthkohli):

Let the smaller number be \(\rm n\). Then what is the expression for the average?

OpenStudy (anonymous):

I'm confuse

Parth (parthkohli):

Do you know what average of two numbers is?

OpenStudy (anonymous):

No ...that's the whole question

OpenStudy (anonymous):

n = the smaller number (35+n)/2 = expression to the average of number Then, we can write this problem to this equation: \[\frac{ 35+n }{ 2 } = 2 (35-n)\] 35+n = 4 (35-n) 35+n = 140 - 4n 4n+n = 140 -35 5n = 105 n=21 You can test the result by entering value of n to the problem again.